OpenStudy (anonymous):

Giving medal for best answer! :3 Solve for f 3(f-q)=18 Would the answer be.. f=18−q over 3?

Nnesha (nnesha):

nope

OpenStudy (anonymous):

Nope, the problem has 18 at the end. :/ It's telling me to solve F.

OpenStudy (anonymous):

and that's not an answer for the problem. :/

OpenStudy (anonymous):

Would it be f=6-q?

Nnesha (nnesha):

is this your question right 3( f - q) = 18 - q right ??????????/

OpenStudy (anonymous):

Nope, just 3(f-q)=18

Nnesha (nnesha):

ohhh ok i'm sorry i thought there is -q on right side because you post your answer so i was confuse with both sorry again

OpenStudy (anonymous):

No, it's completely fine!

Nnesha (nnesha):

ok so divide both side by 3 you should get f - q = 6 right so after that you have to add q both side

Nnesha (nnesha):

 so answer should be what